# Designing Speech Technologies for Australian Aboriginal English: Opportunities, Risks and Participation

### v2

@@ 3. Opportunities and Risks for ASR for Australian Aboriginal English
With the contexts and questions of the previous section in mind, we now consider Automated Speech Recognition (ASR) systems for Aboriginal English, and the potential opportunities and risks. We consider both deployed ASR systems as well as the projects which build ASR systems; our ultimate goal is to build inclusive technologies using inclusive methods.
3.1. System Deployment Opportunities
ASR supports many application domains (Aksënova et al., 2021), and improving ASR for Aboriginal English could provide multiple specific benefits to Aboriginal and Torres Strait Islander people, their communities, and non-Indigenous people who are encountering Aboriginal English. We attempt to distinguish those possible benefits with as much specificity as possible, and in contrast to prior work we adopt a human-centered framing rather than a technology-centered one.
Opportunity 1: Improved dictation for Aboriginal English speakers.
@@ 3.1. System Deployment Opportunities
We note that not all dictation applications are likely to benefit Aboriginal English speakers. For example, ASR is sometimes used in dictation practice by second-language learners, however Aboriginal English is not formally taught or studied as a second-language.
@@
ASR can be initiated by those within the speaker’s community or social network, for example, users of a voicemail service providing automated transcriptions. This can be useful in a variety of contexts in which reading is preferred to listening, including in quiet environments in which silence is preferred, in noisy environments in which hearing is difficult, when privacy is desirable in public environments, and for people with hearing disabilities. An Aboriginal English speaker is often unaware when their voice is being transcribed, and the reader might or might not listen to the voice recording to verify fidelity of the transcription. The reader may be able to accommodate and even correct some ASR errors, based on their familiarity with the speaker and/or context, even without listening to the recording.
@@
When speech and video platforms host content, they often provide closed captions. Such content can be unscripted (typical for Aboriginal English on YouTube, for example, see Appendix B) or scripted (common in TV shows, movies, podcasts and vodcasts, etc—for an analysis of scripted Aboriginal English see (Bednarek, 2021)). These are generated offline and stored as metadata alongside the Aboriginal English voice recording. The platforms may have some access controls, but in the general case both the recordings and the transcriptions may be accessible to potentially anyone online, including listeners and readers with no prior experience with Aboriginal English. Listening to the recordings might not help readers without familiarity with Aboriginal English to correct some ASR errors.
Due to systemic societal biases, some applications of closed captioning may encounter relatively little Aboriginal English compared to Australian English. The automated transcription of Australian parliamentary proceedings (Gollan et al., 2005), for example, is one use case for Aboriginal English ASR that is impacted both by the under-representation of Indigenous people in Australian parliaments and by the social stigmatisation of speaking Aboriginal English in formal contexts.
Opportunity 4: Improved human-machine interactions for Aboriginal English speakers.
ASR is increasingly incorporated into a wide range of computational systems used by Aboriginal English speakers. These include automated voice assistants, automated or semi-automated call centres including for critical banking and government services, web search using voice queries, and audio/video retrieval that indexes speech under transcribed terms. (Voice Query-By-Example technologies (Parada et al., 2009) do not require ASR but constitute a minority of audio retrieval systems.) In many cases, the transcription might not be seen by any human, including the Aboriginal English speaker. Instead, the transcription is typically the input to a Natural Language Understanding component, which outputs some form of semantic representation in forms such as a task frame, a search index entry, or an expanded Information Retrieval query (Efthimiadis, 1996; Jurafsky and Martin, 2008). More accurate transcriptions can lead to more relevant search results, more efficient interactions, and fewer abandoned tasks.
3.2. System Deployment Risks
@@
Risk 1: Inappropriate for the contexts of Aboriginal English.
Researchers should not assume that minoritised language communities have the same wants and needs as mainstream language communities (Bird, 2022), nor that multilingual speech communities require the same technologies for all languages (Lent et al., 2022; Bird and Yibarbuk, 2024). For example, would some Aboriginal English speakers prefer to code-switch into Standardised Australian English when interacting with machines, if it results in greater task efficiency? What about when dictating messages to family and friends? Conversely, would they prefer more support to speak Aboriginal English in contexts which are currently not well supported? There is a shortage of research on the real-world needs of Aboriginal English speakers. Building systems that are neither needed or wanted is not only inefficient—it also undermines the rights of minoritised communities to determine which technologies are built for them.
@@
Like all language varieties, Aboriginal English is not homogeneous, but exhibits regional variation (Dickson, 2019). Furthermore, and notably, Aboriginal English includes both acrolectal and basilectal varieties (§2.1). If the training data for an ASR model does not represent sufficient variation, then the model will fail to recognise the breadth of forms of Aboriginal English. If the training data collection protocols are not designed to be culturally safe, then speakers may code switch into Standardised Australian English, reducing the validity of the data.
Risk 3: Psychological harms for Aboriginal English users when dictating or interacting with systems.
Prior research has found that when ASR systems make errors, some members of minoritised language communities feel that the computer “doesn’t like” how they are speaking (Brewer et al., 2023), or that the technology isn’t built for people like them (Mengesha et al., 2021). Whereas speakers of mainstream language varieties may blame the ASR system for poor performance, speakers of minoritised communities may blame themselves, and experience a drop in self esteem or an increase in self consciousness about their identity (Wenzel et al., 2023). They may feel that their language variety is being judged as illegitimate (Cunningham et al., 2024). If ASR fails to meet expectations for speakers of Aboriginal English, all of these impacts may be reproduced.
Risk 4: Torquing of Aboriginal English sociolinguistic identities.
ASR systems function as classifiers not just of speech, but also, by extension, of language varieties and sociolinguistic identities. As such, ASR systems can be examined through Bowker and Star’s sociological lens of torque and residuals (Bowker and Star, 2000). “Residuals” refers to the “other” categories that are relevant to the individual but missing from the classificatory system, while “torque” refers to the painful twisting and bending that occurs when an individual’s data is forced into an inappropriate classificatory system, such as forcing individuals into racial categories in Apartheid South Africa. Previous work has explored how gender classification technologies produce torque and othering for trans people (Scheuerman et al., 2019), and how medical and biometric IT projects can torque minoritised groups (Avlona and Shklovski, 2024; Singh and Jackson, 2017). We posit that a similar form of torque exists for users of non-standardised language varieties, when ASR systems systematically fail to recognise spoken forms salient to social identities, bending and twisting them into the written forms of mainstream dictionaries of colonising cultures.
@@ 3.3. Project Opportunities
We take the position that language technology projects provide the potential for a broad range of opportunities, especially when the project is conducted not just for the community but also with the community (Cameron et al., 1993).
@@
Opportunity 6: Project produces economic opportunities for Indigenous community.
Data collection projects can provide opportunities for what Sloane et al. (2022) call “participation as work”, with equitable compensation. There are large economic disparities in Australia, with median incomes for Indigenous people around 23% lower than for non-Indigenous (Australian Institute for Health and welfare, 2023), and Indigenous people have higher rates of unemployment for all education levels (Australian Institute of Health and Welfare, 2023b). If equitable compensation and non-extractive processes are successfully adopted, data work can create economic opportunities for community.
@@
Providing opportunities for Indigenous control over project outputs and design decision can have not only consequential benefits (e.g., avoiding data set misuse by third parties in ways that do not accord with Indigenous aspirations), but also constitute a form of procedural justice through greater self-determination.
@@ 3.4. Project Risks
Developing inclusive ASR inclusively requires more than addressing model biases (Feng et al., 2024)— it also requires researchers to integrate participatory practices throughout the project. As such, we take the position that, when developing language technologies, risks need to be considered using a wide-angle lens that takes into account not just the deployment scenarios but also the project processes used when developing the technology. For projects with Indigenous communities, “it’s how do you things that matters” (Cooper et al., 2024), and the ends should not be used to justify the means.
@@
A lack of community participation, generally, can not only lead to poorer data and models, but can also miss opportunities to address social injustices (Sloane et al., 2022). Failure to partner when appropriate with Indigenous communities, specifically, can reproduce histories of ignoring or disrespecting Indigenous sovereignty. Calls for collaborative and participatory research recognise community rights to be meaningfully included in language research that involves them (Carew et al., 2015; Cameron et al., 1993). This is emphasised by unesco in the Los Pinos Declaration calling for a decade of action on Indigenous languages: “Nothing about us without us” (UNESCO, United Nations Educational and Organization).
@@
Recording one’s own voice can be an uncomfortable experience for many people, and especially so for members of minoritised communities encountering researchers from universities or technology companies using mainstream English. Unless care is taken, prompts might not be culturally relevant, verge into territories of secret traditional knowledge, elicit personal narratives of trauma, or trigger feelings that their culture is not important and valid.
@@ 4. Designing ASR for Australian Aboriginal English
We now discuss a real-world partnership between Google and University of Western Australia which aims to improve ASR for Australian Aboriginal English. We focus on the design considerations, and specifically on our efforts to relate the project goals and processes to both the sociolinguistic context (§2.1) and the technological opportunities (§3.1). We also discuss how we attempt to mitigate the system and project risks (§3.2, §3.4) while providing societal and economic opportunities (§3.3), e.g., by integrating participatory processes throughout the project lifecycle.
When building technologies for Indigenous speech communities—of either ancestral or contact varieties—it is important to consider the ways in which projects face social and technical questions distinct from those that involve mainstream languages (see, e.g., Pine et al., 2024; Bird, 2020b; Carew et al., 2015; Noe and Kirshenbaum, 2024; Lent et al., 2022; Le Ferrand et al., 2022; Taylor et al., 2020). For example, ASR development pipelines typically make assumptions about speech communities and language use, including the existence of ‘gold standard’ spoken and written forms (Markl et al., 2023). These assumptions break down in the case of Aboriginal English, for which the written form is not standardised, and so community participation in the design process is a prerequisite for building technologies that serve the community.
@@ 4.1. System desiderata
ASR should support the use of Aboriginal English in all application horizontals which incorporate ASR systems. These include: 1) dictation of messages and documents; 2) voice queries to search engines; 3) voice commands to virtual assistants; and 4) automated transcriptions of voicemail messages, meetings, podcasts, and video content (Aksënova et al., 2021). It is important to distinguish applications in which the transcripts will be processed by computers (Opportunity 4 in §3, e.g., commands to virtual assistants) from those where the transcripts will be read by humans (Opportunities 1–3, e.g., voicemail transcriptions). In the first case, post-processing converts the transcripts to a structured or semi-structured representation, such as an expanded Information Retrieval query or a task frame (Jurafsky and Martin, 2008). This post-processing involves increasing abstraction and standardisation, mapping diverse user inputs into more computationally tractable forms. In the case of transcripts for humans, however, standardisation of linguistic forms is no longer a computational requirement. Instead, standardisation of transcripts minimises the diverse ways in which social meanings are conveyed through linguistic choices. In the process, standardisation reproduces the minoritisation of non-mainstream sociolinguistic identities (Markl et al., 2023). This is especially relevant for informal registers, such as those used in personal messages or social media, in which expressions of social identities are more prevalent and varied.
@@ 4.3. Transcription of speech data
The introduction of writing by settler-colonialists has significantly impacted Indigenous populations (Ong, 2013; Franchetto, 2008). In Aboriginal English, there are two major questions regarding orthographic practices. First, how are loan words from ancestral languages, including place names and Country names (e.g. ‘Nyungar’), written in Latin-based script? Second, how do Indigenous people in Australia deliberately adopt their own spelling conventions, e.g., ‘Blak’. These are not purely technical questions, but also sociopolitical (Franchetto, 2008; Markl et al., 2023; Bucholtz, 2000) (cf. Risk 4).
To co-develop transcription guidelines with community members, we required examples of public Aboriginal English to refer to. For this, we used a corpus of over 100 hours of YouTube videos that are primarily in Aboriginal English (see Appendix B). An Indigenous team, led by Rodríguez Louro, developed the corpus and used their expertise with Aboriginal English to validate the variety spoken, ensuring that speakers were only coded as Aboriginal English if they used at least two linguistic features that distinguish this variety from Standardised Australian English. We developed an initial version of variety-specific transcription guidelines using examples from this corpus. Example guidelines include:
1) borrowings from Pidgin and ancestral languages should be transcribed verbatim, with an English translation included;
2) variety-specific utterance tags should be included, e.g., ‘unna’ and ‘inni’.

### v3

@@ 3. Opportunities and Risks for ASR for Australian Aboriginal English
We now consider Automated Speech Recognition (ASR) systems for Australian Aboriginal English, acknowledging that this focus on converting spoken language to written language itself risks reproducing beliefs about the primary importance of written language vis-à-vis spoken (Chrupała, 2023). With the contexts and questions of the previous section in mind, we consider potential opportunities and risks, informed by the literatures on ASR, Human-computer Interaction, and Indigenous language technologies. We consider both deployed ASR technologies as well as the projects which build ASR technologies; our ultimate goal is to build inclusive technologies using inclusive methods.
3.1. Deployment Opportunities
ASR supports many application domains (Aksënova et al., 2021), and improving ASR for Australian Aboriginal English could provide multiple specific benefits to First Nations people in Australia, their communities, and non-Indigenous people who are encountering Australian Aboriginal English. We attempt to distinguish those possible benefits with as much specificity as possible, and in contrast to prior work we adopt a human-centered framing rather than a technology-centered one.
Opportunity 1: Improved dictation for Australian Aboriginal English speakers.
@@ 3.1. Deployment Opportunities
We note that not all dictation applications are likely to benefit Australian Aboriginal English speakers. For example, ASR is sometimes used in dictation practice by second-language learners, however Australian Aboriginal English is not formally taught or studied as a second-language.
@@
ASR can be initiated by those within the speaker’s community or social network, for example, users of a voicemail service providing automated transcriptions. This can be useful in a variety of contexts in which reading is preferred to listening, including in quiet environments in which silence is preferred, in noisy environments in which hearing is difficult, when privacy is desirable in public environments, and for people with hearing disabilities. An Australian Aboriginal English speaker is often unaware when their voice is being transcribed, and so is unlikely to modify their voice. The reader might or might not listen to the voice recording to verify fidelity of the transcription. The reader may be able to accommodate and even correct some ASR errors, based on their familiarity with the speaker and/or context, even without listening to the recording.
@@
When speech and video platforms host content, they often provide closed captions. Such content can be unscripted (typical for Australian Aboriginal English on YouTube, for example, see Appendix B) or scripted (common in TV shows, movies, podcasts and vodcasts, etc—for an analysis of scripted Australian Aboriginal English see (Bednarek, 2021)). These are generated offline and stored as metadata alongside the Australian Aboriginal English voice recording. The platforms may have some access controls, but in the general case both the recordings and the transcriptions may be accessible to potentially anyone online, including listeners and readers with no prior experience with Australian Aboriginal English. Listening to the recordings might not help readers without familiarity with Australian Aboriginal English to correct some ASR errors.
Due to systemic societal biases, some applications of closed captioning may encounter relatively little Australian Aboriginal English compared to Australian English. The automated transcription of Australian parliamentary proceedings (Gollan et al., 2005), for example, is one use case for Australian Aboriginal English ASR that is impacted both by the under-representation of Indigenous people in Australian parliaments and by the social stigmatisation of speaking Australian Aboriginal English in formal contexts.
Opportunity 4: Improved human-machine interactions for Australian Aboriginal English speakers.
ASR is increasingly incorporated into a wide range of computational systems used by Australian Aboriginal English speakers. These include automated voice assistants, automated or semi-automated call centres including for critical banking and government services, web search using voice queries, and audio/video retrieval that indexes speech under transcribed terms. (Voice Query-By-Example technologies (Parada et al., 2009) do not require ASR but constitute a minority of audio retrieval systems.) In many cases, the transcription might not be seen by any human, including the Australian Aboriginal English speaker. Instead, the transcription is typically the input to a Natural Language Understanding component, which outputs some form of semantic representation in forms such as a task frame, a search index entry, or an expanded Information Retrieval query (Efthimiadis, 1996; Jurafsky and Martin, 2008). More accurate transcriptions can lead to more relevant search results, more efficient interactions, and fewer abandoned tasks.
3.2. Deployment Risks
@@
Risk 1: Inappropriate for the contexts of Australian Aboriginal English.
Researchers should not assume that minoritised language communities have the same wants and needs as mainstream language communities (Bird, 2022), nor that multilingual speech communities require the same technologies for all languages (Lent et al., 2022; Bird and Yibarbuk, 2024). For example, would some Australian Aboriginal English speakers prefer to switch into Standardised Australian English when interacting with machines, if it results in greater task efficiency? What about when dictating messages to family and friends? Conversely, would they prefer more support for speaking Australian Aboriginal English in contexts which are currently not well supported? There is a shortage of research on the real-world needs of Australian Aboriginal English speakers. Building systems that are neither needed or wanted is not only inefficient—it also undermines the rights of minoritised communities to determine which technologies are built for them.
@@
Like all language varieties, Australian Aboriginal English is not homogeneous, but exhibits regional variation (Dickson, 2019). Furthermore, and notably, Australian Aboriginal English includes both acrolectal and basilectal varieties (§2.1). If the training data for an ASR model does not represent sufficient variation, then the model will fail to recognise the breadth of forms of Australian Aboriginal English. If the training data collection protocols are not designed to be culturally safe, then speakers may switch into Standardised Australian English, reducing the validity of the data.
Risk 3: Psychological harms for Australian Aboriginal English users when dictating or interacting with systems.
Prior research has found that when ASR systems make errors, some members of minoritised language communities feel that the computer ‘doesn’t like’ how they are speaking (Brewer et al., 2023), or that the technology isn’t built for people like them (Mengesha et al., 2021). Whereas speakers of mainstream language varieties may blame the ASR system for poor performance, speakers of minoritised communities may blame themselves, and experience a drop in self esteem or an increase in self consciousness about their identity (Wenzel et al., 2023). They may feel that their language variety is being judged as illegitimate (Cunningham et al., 2024). If ASR fails to meet expectations for speakers of Australian Aboriginal English, all of these impacts may be reproduced. Feelings of shame or embarrassment might also arise if speakers are self-conscious about their way of speaking being transcribed and re-contextualised into contexts where mainstream written English is the norm.
Risk 4: Torquing of Australian Aboriginal English sociolinguistic identities.
ASR systems function as classifiers not just of speech, but also, by extension, of language varieties and sociolinguistic identities. As such, ASR systems can be examined through Bowker and Star’s sociological lens of torque and residuals (Bowker and Star, 2000). ‘Residuals’ refers to the ‘other’ categories that are relevant to the individual but missing from the classificatory system, while ‘torque’ refers to the painful twisting and bending that occurs when an individual’s data is forced into an inappropriate classificatory system, such as forcing individuals into racial categories in Apartheid South Africa. Previous work has explored how gender classification technologies produce torque and othering for trans people (Scheuerman et al., 2019), and how medical and biometric IT projects can torque minoritised groups (Avlona and Shklovski, 2024; Singh and Jackson, 2017). We posit that a similar form of torque exists for users of non-standardised language varieties, when ASR systems systematically fail to recognise spoken forms salient to social identities, bending and twisting them into the written forms of mainstream dictionaries of colonising cultures.
@@ 3.3. Project Opportunities
We take the position that language technology projects provide the potential for a range of opportunities, especially when the project is conducted not just for the community but also with the community (Cameron et al., 1993).
@@
Opportunity 6: Project work produces opportunities for Indigenous community.
Data collection projects can provide opportunities for what Sloane et al. (2022) call ‘participation as work’, with equitable compensation. There are large economic disparities in Australia, with median incomes for Indigenous people around 23% lower than for non-Indigenous (Australian Institute for Health and welfare, 2023), and Indigenous people have higher rates of unemployment for all education levels (Australian Institute of Health and Welfare, 2023). If equitable compensation and non-extractive processes are successfully adopted, data work can create economic opportunities for community, as well as opportunities for meaningful cultural engagement (Bird, 2022).
@@
Providing opportunities for Indigenous control over project outputs, for example through data governance. Design decisions can have not only consequential benefits (e.g., avoiding data set misuse by third parties in ways that do not accord with Indigenous aspirations), but also constitute a form of procedural justice through greater self-determination.
@@ 3.4. Project Risks
Developing inclusive ASR inclusively requires more than addressing model biases (Feng et al., 2024)— it also requires researchers to integrate participatory practices throughout the project. As such, we take the position that, when developing language technologies, risks need to be considered using a wide-angle lens that takes into account not just the deployment scenarios but also the project processes used when developing the technology. For projects with Indigenous communities, ‘it’s how do you things that matters’ (Cooper et al., 2024), and the ends should not be used to justify the means.
@@
A lack of community participation, generally, can not only lead to poorer data and models, but can also miss opportunities to address social injustices (Sloane et al., 2022). Failure to partner when appropriate with Indigenous communities, specifically, can reproduce histories of ignoring or disrespecting Indigenous sovereignty. Calls for collaborative and participatory research recognise community rights to be meaningfully included in language research that involves them (Carew et al., 2015; Cameron et al., 1993). This is emphasised by unesco in the Los Pinos Declaration calling for a decade of action on Indigenous languages: ‘Nothing about us without us’ (UNESCO, United Nations Educational and Organization).
@@
Recording one’s own voice can be an uncomfortable experience for many people, and especially so for members of minoritised communities encountering researchers from universities or technology companies using mainstream English. Unless care is taken, prompts might not be culturally relevant, stray into territories of secret traditional knowledge, elicit personal narratives of trauma, or trigger feelings that their culture is not important and valid. Data work may also risk being meaningless to workers’ social or cultural contexts, for example, transcribing long recordings of a non-standardised variety into a mainstream script can be tedious and unengaging work (Bird, 2022, 2021).
@@ 4. Designing ASR for Australian Aboriginal English
We now discuss a real-world partnership between Google and The University of Western Australia which aims to improve ASR for speakers of Australian Aboriginal English. We focus on the design considerations, and specifically on our efforts to relate the project goals and processes to both the sociolinguistic context (§2.1) and the technological opportunities (§3.1). We discuss not just technical questions, but the design of the project itself, including how we attempt to mitigate the system and project risks (§3.2, §3.4) while providing societal and economic opportunities (§3.3), e.g., by integrating participatory processes throughout the project lifecycle.
When building technologies for Indigenous speech communities—of either ancestral or contact varieties—it is important to consider the ways in which projects face social and technical questions distinct from those that involve mainstream languages (see, e.g., Pine et al., 2024; Bird, 2020b; Carew et al., 2015; Noe and Kirshenbaum, 2024; Lent et al., 2022; Le Ferrand et al., 2022; Taylor et al., 2020). For example, ASR development pipelines typically make assumptions about speech communities and language use, including the existence of ‘gold standard’ spoken and written forms (Markl et al., 2023). These assumptions break down in the case of Australian Aboriginal English, for which the written form is not standardised, and so community participation in the design process is a prerequisite for building technologies that serve the community.
@@ 4.1. System desiderata
ASR should support the use of Australian Aboriginal English in all application horizontals which incorporate ASR systems. These include: 1) dictation of messages and documents; 2) voice queries to search engines; 3) voice commands to virtual assistants; and 4) automated transcriptions of voicemail messages, meetings, podcasts, and video content (Aksënova et al., 2021). It is important to distinguish applications in which the transcripts will be processed by computers (Opportunity 4 in §3, e.g., commands to virtual assistants) from those where the transcripts will be read by humans (Opportunities 1–3, e.g., voicemail transcriptions). In the first case, post-processing converts the transcripts to a structured or semi-structured representation, such as an expanded Information Retrieval query or a task frame (Jurafsky and Martin, 2008). This post-processing involves increasing abstraction and standardisation, mapping diverse user inputs into more computationally tractable forms. In the case of transcripts for humans, however, standardisation of linguistic forms is no longer a computational requirement. [^5] The requirement for readability by non-linguists also precludes transcription into phonetic or phonemic representations, as has been done for some other work such as (Michaud et al., 2018). Instead, standardisation of transcripts minimises the diverse ways in which social meanings are conveyed through linguistic choices. In the process, standardisation reproduces the minoritisation of non-mainstream sociolinguistic identities (Markl et al., 2023). This is especially relevant for informal registers, such as those used in personal messages or social media, in which expressions of social identities are more prevalent and varied.
@@ 4.2. Elicitation of speech data
Once collected, all data will be screened by a team of Indigenous RAs, led by Collard, who will vet the audio files and transcriptions for cultural sensitivity. Any culturally unsafe material (e.g., stories of specific people, places, Indigenous knowledges, trauma, and loss) will be erased from the corpus before sharing the data with others.
@@ 4.3. Transcription of speech data
The introduction of writing by settler-colonialists has significantly impacted Indigenous populations (Ong, 2013; Franchetto, 2008). In Australian Aboriginal English, there are two major questions regarding orthographic practices. First, how are loan words from ancestral languages, including place names and Country names (e.g. ‘Nyungar’), written in Latin-based script? Second, how do Indigenous people in Australia deliberately adopt their own spelling conventions, e.g., ‘Blak’. These are not purely technical questions, but also sociopolitical (Franchetto, 2008; Markl et al., 2023; Bucholtz, 2000) (cf. Risk 4).
To co-develop transcription guidelines with community members, we required examples of public Australian Aboriginal English to refer to. For this, we used a corpus of over 100 hours of YouTube videos that are primarily in Australian Aboriginal English (see Appendix B). An Indigenous team, led by Rodríguez Louro, developed the corpus and used their expertise with Australian Aboriginal English to validate the variety spoken, ensuring that speakers were only coded as Australian Aboriginal English if they used at least two linguistic features that distinguish this variety from Standardised Australian English. We developed an initial version of variety-specific transcription guidelines using examples from this corpus. Example guidelines include:
  1. (1)
    borrowings from Pidgin and ancestral languages should be transcribed verbatim, with an English translation included;
  2. (2)
    variety-specific utterance tags should be included, e.g., ‘unna’ and ‘inni’.